# Multiple quantum exceptional, diabolical, and hybrid points in multimode bosonic systems: I. Inherited and genuine singularities

### v1

@@ Abstract
The existence and degeneracies of quantum exceptional, diabolical, and hybrid (i.e., diabolically degenerated exceptional) singularities of simple bosonic systems composed of up to five modes with damping and/or amplification are analyzed. Their dynamics governed by quadratic non-Hermitian Hamiltonians is followed using the Heisenberg-Langevin equations. Conditions for the observation of inherited quantum hybrid points, observed directly in the dynamics of field operators, having up to third-order exceptional and second-order diabolical degeneracies are revealed. Exceptional and diabolical genuine points and their degeneracies observed in the dynamics of second-order field-operator moments are analyzed. Surprisingly, exceptional degeneracies of only second and third orders are revealed. Nevertheless the analyzed bosonic systems exhibit rich dynamics, also owing to their common second-order diabolical degeneracies.
I Introduction
@@
Subsequent studies have revealed also other features observed in $\mathcal{PT}$-symmetric systems in addition to EPs. For example, the spectral degeneracies which are not accompanied by the corresponding eigenvector degeneracies, were observed [42]. Such degeneracies do not usually lead to the above discussed physical effects and so the corresponding points in the system parameter space were named diabolical points (DPs). Note that DPs, contrary to EPs, can also be observed in Hermitian systems [42]. As pointed out in [43], it may happen that this DP or rather diabolical degeneracy (DD) occurs at an EP. In this case, ‘independent’ (i.e. with different eigenvectors) multiple spectral and eigenvector degeneracies are found in systems and we refer to them as hybrid diabolic exceptional points (HPs). Such systems then naturally exhibit a physical behavior similar to that observed at EPs. Moreover, new effects originating in diabolical degeneracy may arise. For example, the system behavior when encircling an HP has been used to construct a multi-mode optical switch [44].
@@
Non-Hermitian $\mathcal{PT}$-symmetric optical bosonic systems are especially interesting from the point of view of their behavior at EPs. Their infinite-dimensional Hilbert space leads to numerous manifestations of modifications of their dynamics at quantum EPs (QEPs), i.e. EPs observed in quantum systems [18], including the effect of quantum jumps [45, 46]. The system dynamics may be followed either directly in the Hilbert space [the Liouville space of statistical operators] or in the complementary space of field-operator moments (FOMs) of all orders [47, 48]. The studies performed in the moment space of field operators in multimode bosonic systems described by non-Hermitian quadratic Hamiltonians revealed different types of degeneracies related to QEPs [49, 43]. Moreover, they allowed to sort QEPs into three classes: inherited QEPs, genuine QEPs and induced QEPs. The dynamical equations for the mean values of field operators indicated the presence of inherited QEPs and quantum HPs (QHPs) [43] that represent the core of the studied unusual behavior. The presence of such inherited QEPs and QHPs then implies the existence of genuine QEPs and QHPs [43] observed in the dynamics of higher-order FOMs. With the increasing FOM order, the degeneracies of genuine QEPs and QHPs increase. Moreover, similar or identical FOMs, as being related by the field commutation relations, arise in the formal construction of higher-order FOMs. Thus, we can also define induced QEPs and QHPs [43] that further enlarge the multiplicity of the spectral degeneracies. However, as these redundant FOMs share their time evolution with the FOMs contributing to genuine QEPs and QHPs, they do not lead to additional diversity of the system evolution. Thus, they are not interesting from the point of view of the dynamics of FOMs of a given order. Thus, this dynamics is fully characterized by the corresponding genuine QEPs and QHPs. As the properties of genuine QEPs and QHPs originate in those of the inherited QEPs and QHPs, the analysis of the latter is crucial for the understanding of a system evolution. For this reason, it is important to identify the inherited QEPs and QHPs and their degeneracies in simple bosonic systems formed by smaller numbers of bosonic modes. This analysis may then be exploited for further studies of physical effects in such systems.
@@
The system composed of two mutually interacting modes, one being damped and the other amplified, was already analyzed from this point of view in Ref. [43]. This analysis may be considered as the simplest building block useful for investigations of more complex bosonic systems. Here, we consider systems composed of up to five modes in different configurations promising for the observation of QEPs and QHPs. Looking for inherited QEPs with higher-order EDs is the main goal of our investigations. It is motivated by the fact that the higher is the ED order, the more modified is the system dynamics at a QEP. This then enhances the physical effects specific to QEPs like improvement in measurement precision or enhancement of nonlinear effects.
@@
As we have been able to reveal only up to the third-order ED of QEPs of the analyzed systems, we continue our analysis in Ref. [50], being the second part of this paper, in which we pay attention to the spectral degeneracies of non-Hermitian Hamiltonians observed only in specific subspaces of the systems’ Liouville spaces as well as non-Hermitian Hamiltonians with unidirectional coupling. In Ref. [50], we also address numerical identification of QEPs and QHPs. Moreover, we extend the analysis of the genuine and induced QEPs and QHPs by considering the FOMs of a general order.
@@
The paper is organized as follows. A two-mode bosonic system with unequal damping and/or amplification rates, as the simplest considered model, is analyzed in Sec. II. Section III brings the analysis of a related three-mode linear system. The corresponding generalized four-mode linear and circular systems are investigated in Sec. IV whereas the analysis of the five-mode linear and pyramid systems is found in Sec. V. Section VI brings conclusions.
II Two-mode bosonic system: Basic building blocks
@@
We begin with the consideration of one of the simplest $\mathcal{PT}$-symmetric bosonic systems that is composed of two modes: one being attenuated and the other amplified [for the scheme, see Fig. 1(a)]. We note that even a one-mode bosonic system may exhibit $\mathcal{PT}$-symmetric behavior, as shown in Ref. [51]. We also pay attention only to the systems described by quadratic Hamiltonians that lead to linear exactly-solvable Heisenberg equations. Though these Hamiltonians lead to linear dynamical equations, they allow for describing the nonlinear effect of photon-pair generation and annihilation. This effect is commonly used in various quantum optical systems to generate entangled [52] and squeezed [53, 54] states of light. $\mathcal{PT}$-symmetry restricts the form of the studied non-Hermitian Hamiltonian such that the underlying dynamics is described by the dynamical matrix composed of typical $2\times 2$ submatrices. They are also used to build the matrices of more complex $\mathcal{PT}$-symmetric bosonic systems. Moreover, we also consider bosonic systems in which damping and amplification are not in balance, which is a typical situation of $\mathcal{PT}$-symmetric systems. In this unbalanced case, relying on the results presented in Ref. [55], we may introduce a specific interaction picture in which the average damping or amplification dynamics is projected out and the remaining dynamics exhibits the features found in $\mathcal{PT}$-symmetric systems.
@@
Figure 1: Schematic diagrams of the bosonic systems composed of (a) two, (b) three, (c—e) four, and (f,g) five modes with typical linear, circular, and pyramid configurations that exhibit QHPs. Strengths $\epsilon$ and $\kappa$ characterize, respectively, the linear and nonlinear coupling between the modes, $\gamma$, with subscripts indicating the mode number(s), are the damping or amplification rates, and annihilation operators $\hat{a}$ identify the mode number via their subscripts, and $\gamma_{jk}$ indicates that $\gamma_{j}=\gamma_{k}$.
@@
[…]
@@
is expressed in terms of $2\times 2$ submatrices $\bm{\tilde{\gamma}_{j}}$, $j=1,2$, and $\bm{\xi}$ defined as:
@@
The $2\times 2$ matrices, $\bm{\tilde{\gamma}_{j}}$ ($j=1,2$) and $\bm{\xi}$, given in Eq. (10) can be simultaneously diagonalized using the diagonalization transformation appropriate to the matrix $\bm{\xi}$ as the remaining two matrices are linearly proportional to the unity matrix and, thus, are not modified by the transformation. This diagonalization transformation then factorizes (decomposes) the matrix $\bm{M^{(2)}}$ in Eq. (5) into two independent $2\times 2$ matrices with the structure of the original matrix $\bm{M^{(2)}}$.
Denoting an eigenvalue of the matrix $\bm{\xi}$ as $\xi$, we may express the eigenfrequencies $\lambda_{1,2}^{M^{(2)}}$ and eigenvectors $\bm{y_{1,2}^{M^{(2)}}}$ of the factorizing (decomposing) $2\times 2$ matrices of the matrix $\bm{M^{(2)}}$ in Eq. (2) in the following common form:
@@
and $\beta=\sqrt{\zeta^{2}-\gamma_{-}^{2}}$. We note that the average damping or amplification rate $\gamma_{+}=0$ in the usual $\mathcal{PT}$-symmetric systems.
@@
$\bm{Y_{3,4}^{M^{(2)}}}$ $=$ [(±$\left[\begin{array}[]{cccccc}\frac{(\pm\epsilon-\zeta)\chi^{*}}{\kappa\zeta},&\mp\frac{\chi^{*}}{\zeta},&-\frac{(\epsilon\mp\zeta)}{\kappa},&1\end{array}\right]^{T},$ (20)
where $\chi=i\gamma_{-}+\beta$.
@@
These findings about the structure of the matrices describing the dynamical equations have their counterpart in the structure of the analyzed two-mode $\mathcal{PT}$-symmetric system. The $2\times 2$ matrix $\bm{\xi}$, defined in Eq. (10), connects pairs of the annihilation and creation operators of different modes. As such it forms the basic building block of more complex $\mathcal{PT}$-symmetric systems, together with the damping and/or amplification matrices $\bm{\tilde{\gamma}_{j}}$ in Eq. (10). In more complex $\mathcal{PT}$-symmetric systems, the dependencies of the eigenvalues of the dynamical matrices $\bm{M^{(n)}}$ ($n=2,3,\ldots$) on the eigenvalues $\lambda^{\xi}_{1,2}$ of the $\bm{\xi}$ matrix are typically quadratic. This, thus, results in the observation of second-order DD in the dynamical features of the matrices $\bm{M^{(n)}}$. In this case, the eigenvalue analysis of the considered systems considerably simplifies and we may restrict our attention to only the $n\times n$ matrices $\bm{M^{(n)}}$, when a detailed eigenvalue analysis is performed.
The condition ${\beta}=0$ for observing QHPs can be analyzed in the space of system parameters $(\epsilon,\kappa,\gamma_{1},\gamma_{2})$ as follows. The quadratic Hamiltonian $\hat{H}^{(2)}$ in Eq. (1) provides the linear Heisenberg-Langevin equations that allow for the temporal rescaling $\epsilon t$, i.e., only the relative parameters $(\kappa/\epsilon,\gamma_{1}/\epsilon,\gamma_{2}/\epsilon)$ suffice in characterizing the system dynamics. Moreover, the structure of the analyzed systems is such that the average damping or amplification rates $\gamma_{+}$ influences equally only the eigenvalues, but it does not modify the eigenvectors. This makes the space of the system parameters effectively two-dimensional with the spanning parameters $(\kappa/\epsilon,\gamma_{-}/\epsilon)$. Using these parameters, the condition $\beta=0$ is expressed as follows:
@@
Thus, the QHPs form a circle with the unit radius in the space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$, as shown in Fig. 1. We note that, at this circle, there is a specific point at $\gamma_{-}=0$ in which all four eigenvectors, given in Eq. (16), are the same which gives raise to the fourth-order QEP. However, this corresponds to the system in which both modes are equally damped or amplified. It is worth noting that the eigenvectors of both matrices $\bm{M^{(2)}}$, given in Eq. (5), and $\bm{\xi}$, in Eq. (10), are degenerated at this point.
@@
Figure 2: Real parts $\lambda^{\rm r}$ of the eigenvalues (a) $\lambda_{1,2}^{M^{(2)}}$ of the matrix $\bm{M^{(2)}}$, given in Eq. (15), for the two-mode bosonic system, (b) $\lambda_{1,2,3}^{M^{(3)}}$ of the matrix $\bm{M^{(3)}}$, given in Eq. (35), for the three-mode linear bosonic system, (c) [(d)] $\lambda_{1,\ldots,4}^{M^{(4)}_{\rm l1}}$ [$\lambda_{1,\ldots,4}^{M^{(4)}_{\rm l2}}$] of the matrix $\bm{M^{(4)}_{\rm l1}}$ [$\bm{M^{(4)}_{\rm l2}}$], given in Eqs. (44), (45), and [(49)] for the four-mode linear bosonic system with neighbor modes having equal [different] damping and/or amplification rates, (e) $\lambda_{1,\ldots,4}^{M^{(4)}_{\rm c1}}$ of the matrix $\bm{M^{(4)}_{\rm c1}}$, given in Eq. (58), for the four-mode circular bosonic system with neighbor modes having equal the damping and/or amplification rates, (f) $\lambda_{2,\ldots,5}^{M^{(5)}_{\rm l}}$ of the matrix $\bm{M^{(5)}_{\rm l}}$, given by Eq. (69), for the five-mode linear bosonic system, and (g,h) $\lambda_{2,\ldots,5}^{M^{(5)}_{\rm p}}$ of the matrix $\bm{M^{(5)}_{\rm p}}$, given in Eq. (81), for the five-mode pyramid bosonic system assuming (g) $\beta_{1}=0$ and (h) $\beta_{2}=0$. The eigenvalues are drawn in the parameter space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$, where $\epsilon$ ($\kappa$) is the linear (nonlinear) coupling strength and $\gamma_{-}$ the difference of the damping/amplification rates in individual models. Dashed red curves indicate the positions of the QHPs given by (a,e,g) Eq. (22), (b) Eq. (37), (c) Eq. (48), (d) Eq. (52), (f) Eq. (74), and (h) Eq. (86).

### v3

@@ Abstract
The existence and degeneracies of quantum exceptional, diabolical, and hybrid (i.e., diabolically degenerated exceptional) singularities of simple bosonic systems composed of up to five modes with damping and/or amplification are analyzed. Their dynamics governed by quadratic non-Hermitian Hamiltonians is followed using the Heisenberg-Langevin equations. Their dynamical matrices generally exhibit specific structures that allow for an effective reduction of their dimension by half. This facilitates analytical treatment and enables efficient spectral analysis based on characteristic second-order diabolical degeneracies. Conditions for the observation of inherited quantum hybrid points, observed directly in the dynamics of field operators, having up to third-order exceptional and second-order diabolical degeneracies are revealed. Surprisingly, exceptional degeneracies of only second and third orders are revealed, even though the systems with up to five modes are considered. Exceptional and diabolical genuine points and their degeneracies observed in the dynamics of second-order field-operator moments are also analyzed. Each analyzed bosonic system exhibits its own unique and complex dynamical behavior.
1 Introduction
@@
Subsequent studies have revealed also other features observed in $\mathcal{PT}$-symmetric systems in addition to EPs. For example, the spectral degeneracies which are not accompanied by the corresponding eigenvector degeneracies, were observed [42]. Such degeneracies do not usually lead to the above discussed physical effects and so the corresponding points in the system parameter space were named diabolical points (DPs). Note that DPs, contrary to EPs, can also be observed in Hermitian systems [42]. As pointed out in [43], it may happen that this DP with its diabolical degeneracy (DD) occurs at an EP. In this case, ‘independent’ (i.e. with different eigenvectors) multiple spectral and eigenvector degeneracies are found in systems and we refer to them as hybrid diabolic exceptional points (HPs). Such systems then naturally exhibit a physical behavior similar to that observed at EPs. Moreover, new effects originating in diabolical degeneracy may arise. For example, the system behavior when encircling an HP has been used to construct a multi-mode optical switch [44].
@@
Non-Hermitian $\mathcal{PT}$-symmetric optical bosonic systems are especially interesting from the point of view of their behavior at EPs. Their infinite-dimensional Hilbert space leads to numerous manifestations of modifications of their dynamics at quantum EPs (QEPs), i.e. EPs observed in quantum systems [18], including the effect of quantum jumps [45, 46]. The system dynamics may be followed either directly in the Hilbert space [the Liouville space of statistical operators] or in the complementary space of field-operator moments (FOMs) of all orders [47, 48]. The studies performed in the moment space of field operators in multimode bosonic systems described by non-Hermitian quadratic Hamiltonians revealed different types of degeneracies related to QEPs [49, 43]. Moreover, they allowed to sort QEPs into three classes: inherited QEPs, genuine QEPs and induced QEPs. The dynamical equations for the mean values of field operators indicated the presence of inherited QEPs and quantum HPs (QHPs) [43] that represent the core of the studied unusual behavior. The presence of such inherited QEPs and QHPs then implies the existence of genuine QEPs and QHPs [43] observed in the dynamics of higher-order FOMs. With the increasing FOM order, the degeneracies of genuine QEPs and QHPs increase. Moreover, similar or identical FOMs, as being related by the field commutation relations, arise in the formal construction of higher-order FOMs. Thus, we can also define induced QEPs and QHPs [43] that further enlarge the multiplicity of the spectral degeneracies. However, as these redundant FOMs share their time evolution with the FOMs contributing to genuine QEPs and QHPs, they do not lead to additional diversity of the system evolution. Thus, they are not interesting from the point of view of the dynamics of FOMs of a given order. This dynamics is fully characterized by the corresponding genuine QEPs and QHPs. As the properties of genuine QEPs and QHPs originate in those of the inherited QEPs and QHPs, the analysis of the latter is crucial for the understanding of a system evolution. For this reason, it is important to identify the inherited QEPs and QHPs and their degeneracies in simple bosonic systems formed by smaller numbers of bosonic modes. This analysis may then be exploited for further studies of physical effects in such systems.
@@
The system composed of two mutually interacting modes, one being damped and the other amplified, was already analyzed from this point of view in Ref. [43]. This analysis may be considered as the simplest building block useful for investigations of more complex bosonic systems in which much richer and diverse physical behavior is expected. Here, we consider systems composed of up to five modes in different configurations promising for the observation of QEPs and QHPs. Looking for inherited QEPs with higher-order EDs is the main goal of our investigations. It is motivated by the fact that the higher is the ED order, the more modified is the system dynamics at a QEP. This then enhances the physical effects specific to QEPs like improvement in measurement precision or enhancement of nonlinear effects.
@@
As, surprisingly, we have been able to reveal only up to the third-order ED of QEPs in the analyzed systems, we continue our analysis in Ref. [50], being the second part of this paper, in which we pay attention to the spectral degeneracies of non-Hermitian Hamiltonians observed only in specific subspaces of the systems’ Liouville spaces as well as non-Hermitian Hamiltonians with unidirectional coupling. In particular, systems with unidirectional coupling exhibit significantly stronger non-Hermitian features compared to those analyzed here, enabling the observation of QEPs and QHPs of arbitrary order. In Ref. [50], we also address numerical identification of QEPs and QHPs, useful in the bosonic systems with higher dimensions. We also extend the analysis of the genuine and induced QEPs and QHPs by considering the FOMs of a general order.
At the end of Introduction, we would like to note that while the existence of higher-order EPs has been discussed previously in the literature, those analyses typically rely on semiclassical models that neglect quantum fluctuations, that are necessary for quantum consistency of the models. Moreover, many prior works investigate the spectra of semiclassical non-Hermitian Hamiltonians, which exhibit infinite (though countable) sets of eigenfrequencies. In contrast, the physically relevant spectra are those of the corresponding dynamical matrices governing the evolution equations (e.g., the Heisenberg ones in the Heisenberg picture), which involve only a small number of frequencies. It is precisely these frequencies that determine the observable behavior of the system. In semiclassical models, higher-order EPs can often be identified with relative ease. However, the validity of such models is generally restricted to specific conditions, such as short evolution times or weak damping/amplification. This raises a critical and largely unexplored question: What are the spectral degeneracies in PT-symmetric bosonic systems under general conditions, where fluctuating quantum forces - governed by fluctuation-dissipation theorems - play a non-negligible role?
@@
The paper is organized as follows. A two-mode bosonic system with unequal damping and/or amplification rates, as the simplest considered model, is analyzed in Sec. II. Section III brings the analysis of a related three-mode system in the linear configuration. The corresponding generalized four-mode systems in their linear and circular configurations are investigated in Sec. IV whereas the analysis of the five-mode systems in their linear and pyramid configurations is found in Sec. V. Section VI brings conclusions. In Appendix A, the eigenvalues and eigenvectors of the general $2n\times 2n$ dynamical matrices are summarized. The structure of the second-order FOMs for the analyzed systems is detailed in the tables provided in Appendix B.
2 Two-mode bosonic system: Basic building blocks
@@
We begin with the consideration of one of the simplest $\mathcal{PT}$-symmetric bosonic systems that is composed of two modes: one being attenuated and the other amplified [for the scheme, see Fig. 1(a)]. We note that even a one-mode bosonic system may exhibit $\mathcal{PT}$-symmetric behavior, as shown in Ref. [51]. We also pay attention only to the systems described by quadratic Hamiltonians that lead to linear exactly-solvable Heisenberg equations. Though these Hamiltonians lead to linear dynamical equations, they allow for describing the nonlinear effect of photon-pair generation and annihilation. This effect is commonly used in various quantum optical systems to generate entangled [52] and squeezed [53, 54] states of light. $\mathcal{PT}$-symmetry restricts the form of the studied non-Hermitian Hamiltonian such that the underlying dynamics is described by the dynamical matrix composed of characteristic $2\times 2$ submatrices. They are also used to build the matrices of more complex $\mathcal{PT}$-symmetric bosonic systems. The characteristic block structure — typical of bosonic systems with quadratic Hamiltonians — forms a critical ingredient of the analytical framework presented below, allowing us to uncover the complex structures of eigenvalues and their associated eigenvectors of the corresponding dynamical matrices in the space of system’s parameters. Moreover, we also consider bosonic systems in which damping and amplification are not in balance, which is a typical situation of $\mathcal{PT}$-symmetric systems. In this unbalanced case, relying on the results presented in Ref. [55], we may introduce a specific interaction picture in which the average damping or amplification dynamics is projected out and the remaining dynamics exhibits the features found in $\mathcal{PT}$-symmetric systems.
@@
Figure 1: Schematic diagrams of the bosonic systems composed of (a) two, (b) three, (c—e) four, and (f,g) five modes with typical linear, circular, and pyramid configurations that exhibit quantum exceptional points (QEPs) and quantum hybrid points (QHPs) with various exceptional degeneracies (EDs) and diabolical degeneracies ($D$s) observed in the dynamics of field-operator moments (FOMs) of different orders. Strengths $\epsilon$ and $\kappa$ characterize, respectively, the linear and nonlinear coupling between the modes, $\gamma$, with subscripts indicating the mode number(s), are the damping or amplification rates, and annihilation operators $\hat{a}$ identify the mode number via their subscripts, and $\gamma_{jk}$ indicates that $\gamma_{j}=\gamma_{k}$.
@@
$\hat{H}_{2}$ $=$ $\left[\hbar\epsilon\hat{a}_{1}^{\dagger}\hat{a}_{2}+\hbar\kappa\hat{a}_{1}\hat{a}_{2}\right]+\textrm{H.c.},$ (1)
@@
is expressed in terms of $2\times 2$ submatrices $\bm{\tilde{\gamma}_{j}}$, $j=1,2$, and $\bm{\xi}$ is defined as:
@@
The $2\times 2$ matrices, $\bm{\tilde{\gamma}_{j}}$ ($j=1,2$) and $\bm{\xi}$, given in Eq. (10) can be simultaneously diagonalized using the diagonalization transformation appropriate to the matrix $\bm{\xi}$ as the remaining two matrices are linearly proportional to the unity matrix and, thus, are not modified by the transformation. This diagonalization transformation then decomposes the $4\times 4$ matrix $\bm{M^{(2)}}$ in Eq. (5) into the direct sum of two independent $2\times 2$ matrices belonging to the eigenvalues $\lambda^{\xi}_{1}$ and $\lambda^{\xi}_{2}$ of the matrix $\bm{\xi}$ [see Eq. (18) below] and having the structure of the original $2\times 2$ matrix $\bm{M^{(2)}}$:
$\bm{M^{(2)}}=\left[\begin{array}[]{cccccccc}{\color[rgb]{0,0,1}-i{\gamma}_{1}/2}&0&{\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0\\ 0&{\color[rgb]{0,1,0}-i{\gamma}_{1}/2}&0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}\\ {\color[rgb]{0,0,1}\lambda^{\xi}_{1}}&0&{\color[rgb]{0,0,1}-i{\gamma}_{2}/2}&0\\ 0&{\color[rgb]{0,1,0}\lambda^{\xi}_{2}}&0&{\color[rgb]{0,1,0}-i{\gamma}_{2}/2}\end{array}\right].$ (15)
This step is critical to our analytical approach, and it is applied analogously to systems with a larger number of bosonic modes. It enables an effective reduction in the dimensionality of the diagonalization problem, thereby allowing the derivation of comprehensive analytical expressions essential for analyzing QEPs and QDPs. Denoting a general eigenvalue of the $2\times 2$ matrix $\bm{\xi}$ as $\xi$, we may express the eigenfrequencies $\lambda_{1,2}^{M^{(2)}}$ and eigenvectors $\bm{y_{1,2}^{M^{(2)}}}$ of the decomposing $2\times 2$ matrices of the $4\times 4$ matrix $\bm{M^{(2)}}$ in Eq. (2) in the following common form:
@@
and $\beta=\sqrt{\zeta^{2}-\gamma_{-}^{2}}$. We note that, in general, we use capital Greek letters $\Lambda$ to denote the eigenvalues of the dynamical matrices $\bm{M}$} in their full dimensions, and lowercase Greek letters $\lambda$ for the eigenvalues in the reduced (half) dimensions. We also note that the average damping or amplification rate $\gamma_{+}=0$ in the usual $\mathcal{PT}$-symmetric systems, that are, however, only specific cases in our general calculations.
@@
$\bm{Y_{3,4}^{M^{(2)}}}$ $=$ $\left[\begin{array}[]{cccccc}\frac{(\pm\epsilon-\zeta)\chi^{*}}{\kappa\zeta},&\mp\frac{\chi^{*}}{\zeta},&-\frac{(\epsilon\mp\zeta)}{\kappa},&1\end{array}\right]^{T},$ (25)
where $\chi=i\gamma_{-}+\beta$. The structure of eigenvalues and eigenvectors of a general dynamical $2n\times 2n$ matrix $\bm{M^{(n)}}$ formed by the eigenvalues $\lambda_{k}^{M^{(n)}}$ and eigenvectors $\bm{y_{k}^{M^{(n)}}}$ of the $n\times n$ matrix $\bm{M^{(n)}}$ expressed using $2\times 2$ submatrices and the eigenvalues $\lambda^{\xi}_{1,2}$ and eigenvectors of the submatrix $\bm{\xi}$ is shown in Fig. 2.
Figure 2: Schematic diagram of the structure of eigenvalues $\Lambda_{j}^{M^{(n)}}$ and the corresponding eigenvectors $\bm{Y_{j}^{M^{(n)}}}$, $j=1,\ldots,2n$, of a general dynamical $2n\times 2n$ matrix $\bm{M^{(n)}}$ built from the eigenvalues $\lambda_{k}^{M^{(n)}}$ and the corresponding eigenvectors $\bm{y_{k}^{M^{(n)}}}$, $k=1,\ldots,n$, of the matrix $\bm{M^{(n)}}$ considered as an $n\times n$ matrix composed of $2\times 2$ submatrices at the positions of its elements; $n=2,3,\ldots$. The eigenvalues $\lambda^{\xi}_{1,2}$ and the corresponding eigenvectors $\bm{y^{\xi}_{1,2}}$ belong to the submatrix $\bm{\xi}$.
@@
These findings about the structure of the matrices describing the dynamical equations have their counterpart in the structure of the analyzed two-mode $\mathcal{PT}$-symmetric system. The $2\times 2$ matrix $\bm{\xi}$, defined in Eq. (10), connects pairs of the annihilation and creation operators of different modes. As such it forms the basic building block of more complex $\mathcal{PT}$-symmetric systems, together with the damping and/or amplification matrices $\bm{\tilde{\gamma}_{j}}$ in Eq. (10). In more complex $\mathcal{PT}$-symmetric systems, the dependencies of the eigenvalues of the dynamical $n\times n$ matrices $\bm{M^{(n)}}$ ($n=2,3,\ldots$) on the eigenvalues $\lambda^{\xi}_{1,2}$ of the $\bm{\xi}$ matrix are typically quadratic. This, thus, results in the observation of second-order DD in the dynamical features of the matrices $\bm{M^{(n)}}$. In this case, the eigenvalue analysis of the considered systems considerably simplifies and we may restrict our attention to only the $n\times n$ matrices $\bm{M^{(n)}}$ with their elements in the form of $2\times 2$ submatrices, when a detailed eigenvalue analysis is performed. The consequences of the eigenvalue analysis are then combined with the above second-order DD.
The condition ${\beta}=0$ for observing QHPs can be analyzed in the space of system parameters $(\epsilon,\kappa,\gamma_{1},\gamma_{2})$ as follows. The quadratic Hamiltonian $\hat{H}^{(2)}$ in Eq. (1) provides the linear Heisenberg-Langevin equations that allow for the temporal rescaling $\epsilon t$, i.e., only the relative parameters $(\kappa/\epsilon,\gamma_{1}/\epsilon,\gamma_{2}/\epsilon)$ suffice in characterizing the system dynamics. Moreover, the structure of the analyzed systems is such that the average damping or amplification rate $\gamma_{+}$ influences equally only the eigenvalues, but it does not modify the eigenvectors. This makes the space of the system parameters effectively two-dimensional with the spanning parameters $(\kappa/\epsilon,\gamma_{-}/\epsilon)$. Using these parameters, the condition $\beta=0$ is expressed as follows:
@@
Thus, the QHPs form a circle with the unit radius in the space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$, as shown by the red dashed curve in Fig. 3(a). We note that, at this circle, there is a specific point at $\gamma_{-}=0$ in which all four eigenvectors, given in Eq. (21), are the same which gives raise to the fourth-order QEP. However, this corresponds to the system in which both modes are equally damped or amplified. It is worth noting that the eigenvectors of both matrices $\bm{M^{(2)}}$, given in Eq. (5), and $\bm{\xi}$, in Eq. (10), are degenerated at this point.
@@
Figure 3: Real parts $\lambda^{\rm r}$ of the eigenvalues (a) $\lambda_{1,2}^{M^{(2)}}$ of the matrix $\bm{M^{(2)}}$, given in Eq. (16), for the two-mode bosonic system, (b) $\lambda_{1,2,3}^{M^{(3)}}$ of the matrix $\bm{M^{(3)}}$, given in Eq. (40), for the three-mode linear bosonic system, (c) [(d)] $\lambda_{1,\ldots,4}^{M^{(4)}_{\rm l1}}$ $\lambda_{1,\ldots,4}^{M^{(4)}_{\rm l2}}$] of the matrix $\bm{M^{(4)}_{\rm l1}}$ $\bm{M^{(4)}_{\rm l2}}$], given in Eq. (51) [(55)] for the four-mode linear bosonic system with neighbor modes having equal [different] damping and/or amplification rates, (e) $\lambda_{1,\ldots,4}^{M^{(4)}_{\rm c1}}$ of the matrix $\bm{M^{(4)}_{\rm c1}}$, given in Eq. (65), for the four-mode circular bosonic system with neighbor modes having equal damping and/or amplification rates, (f) $\lambda_{2,\ldots,5}^{M^{(5)}_{\rm l}}$ of the matrix $\bm{M^{(5)}_{\rm l}}$, given by Eq. (77), for the five-mode linear bosonic system, and (g,h) $\lambda_{2,\ldots,5}^{M^{(5)}_{\rm p}}$ of the matrix $\bm{M^{(5)}_{\rm p}}$, given in Eq. (89), for the five-mode pyramid bosonic system assuming (g) $\beta_{1}=0$ and (h) $\beta_{2}=0$. The eigenvalues are drawn in the parameter space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$, where $\epsilon$ ($\kappa$) is the linear (nonlinear) coupling strength and $\gamma_{-}$ the difference of the damping/amplification rates in individual models. Dashed red curves indicate the positions of the QHPs given by (a,e,g) Eq. (27), (b) Eq. (42), (c) Eq. (53), (d) Eq. (57), (f) Eq. (79), and (h) Eq. (91).